Find the volume.
One side is 8ft., one side is 8 ft., one side is 4 ft., and the width is 2ft.
Right triangular prism
Round your answer to the nearest hundredth, if necessary
I am wondering if I can just multiply 8x8x2 and then divide the answer by 2? I am stuck on this one. Please help.Thanks

Answer by Alan3354(21583)

Find the volume.
One side is 8ft., one side is 8 ft., one side is 4 ft., and the width is 2ft.
Right triangular prism
Round your answer to the nearest hundredth, if necessary
I am wondering if I can just multiply 8x8x2 and then divide the answer by 2?
--------------
Volume = area*width
Find the area of the triangle.
base = 4 ft
h = sqrt(8^2 - 4^2)
h = sqrt(48) = 4sqrt(3)
Area = bh/2
Area = 8sqrt(3)
----------
Vol = Area*2
Vol = 16sqrt(3)